OpenStudy (blackops2luvr):

What is the volume of a cone with a radius of 3 feet and a height of 6 feet? Use 3.14 for pi. Round your answer to the nearest hundredth.

OpenStudy (blackops2luvr):

@iGreen

OpenStudy (campbell_st):

the formula is \[V = \frac{1}{3} \times 3.14 \times r^2 \times h\] you have r = 3 and h = 6 so just substitute and calculate the answer to 2 decimal places

OpenStudy (blackops2luvr):

@igreen

OpenStudy (igreen):

\(V = \dfrac{1}{3} \times 3.14 \times 9 \times 6\) Multiply those. @blackops2luvr

OpenStudy (blackops2luvr):

okay.

OpenStudy (blackops2luvr):

56.52

OpenStudy (blackops2luvr):

100pc.

OpenStudy (igreen):

Yep, 56.52 is correct.

OpenStudy (blackops2luvr):

alright. one more assessment then im done for the day :D